OpenStudy (anonymous):

What number should both sides of the following equation be multiplied by to solve for x? X/6.5= 4 5 6 6.5 26

Parth (parthkohli):

As I said, we have to do the opposite operation already done on the variable side. We are dividing x by 6.5, so isolating x would involve the multiplication of 6.5 to both sides.

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

thanks i so bad at math :)

Parth (parthkohli):

Hmm

Parth (parthkohli):

That'd result into x = 6.5 * 4 x = 13 * 2 x = 26

OpenStudy (anonymous):